# Enhancing Weakly Supervised 3D Medical Image Segmentation through Probabilistic-aware Learning

### v1

@@ 6 Acknowledgment
This work was supported in part by U.S. NIH grants R01GM134020 and P41GM103712, NSF grants DBI-1949629, DBI-2238093, IIS-2007595, IIS-2211597, and MCB-2205148. This work was supported in part by Oracle Cloud credits and related resources provided by Oracle for Research, and the computational resources support from AMD HPC Fund.
Supplementary Material
Appendix 0.A The Algorithm of Target Generation
[…]
Appendix 0.B Sampling of the Dependency Scores
@@
During inference, to approximate the integral of Eq. 9, we sample all the dependency scores independently for $M$ times and calculate the final segmentation output $y^{*}$ where $\alpha_{m}$ denotes the dependency scores sampled at the $m$th time:
@@
The proof of Eq. 9 is established as: Given that the dependency scores within the same PMSA layer are independent of each other, and the dependency scores of deeper PMSA layer are dependent on those of former PMSA layers, as indicated by Eq. 17, Eq. 9 could be written as:
@@
Appendix 0.C The Details of Implementation
Our model inherits the contracting-expanding pattern of UNETR [14] but substitutes the encoder by a stack of Probabilistic Transformer blocks, each connected to the decoder following skip connections. We implemented our method using the PyTorch [28] framework and MONAI [^1] https://monai.io/. All experiments are conducted on a single NVIDIA RTXA5000 GPU with 24GB GPU memory. We set the number of transformer encoders to 12 (L=12) with an embedding size of 768 (K=768). Each patch has a resolution of 16x16x16. During training, we use the AdamW optimizer with an initial learning rate of 0.0001 and a batch size of 1. The number of training iterations was set to 6,000. For inference, we employ a sliding window approach with a 50% overlap. The number of sampled points for different labels is proportional to the volume of the corresponding organ: 200 points for the spleen, 400 points for the liver, and 50 points for each of the right and left kidneys.
Appendix 0.D Training and Inference Strategy
Binary Classification. Unlike binary segmentation map which is common in most deep learning tasks, the proposed probability-based pseudo label suffers from large data size and a single point has confidence scores for multiple organ classes, the sum of which might be greater than 1, which could be ambiguous. Thus, our model is trained to make inference of a single organ class, which is formulated as a binary classification task for each point.
Sampling of Dependency Scores. During training, to accelerate the training process, we sample the dependency scores for only one time, while during inference, the dependency scores are sampled for $M$ times, and the final output is calculated as Eq. 10.
Appendix 0.E Additional Qualitative Results
In Fig. 6 and Fig. 6, we present a comparison of our approach with several fully supervised methods and additionally display some visual results to illustrate that our method attains performance comparable to that of fully supervised ones. The visual comparisons provide compelling evidence of the effectiveness of our method in accurately segmenting the desired regions of interest. Despite being trained with limited supervision, our approach demonstrates competitive performance, highlighting its potential as a viable alternative to fully supervised methods.
Figure 5: Qualitative comparison with fully supervised methods.
Figure 5: Qualitative comparison with fully supervised methods.
Figure 6: Qualitative results of segmentation prediction.
Appendix 0.F Exploring Key Parameters
0.F.1 Impact of the Number of Sampled Points
Table 5: Illustration of the number of sampled points.
Dataset n Dice Score↑ HD95↓
Spleen Liver Left Kidney Right Kidney Spleen Liver Left Kidney Right Kidney
BTCV 50 0.6392 0.1081 0.6276 0.5678 336.25 127.16 135.88 292.41
100 0.8001 0.7307 0.6366 0.4772 174.41 295.54 197.66 312.45
150 0.5462 0.7164 0.3856 0.5756 348.11 304.69 333.5 183.65
200 0.8279 0.8157 0.7599 0.7164 63.08 265.79 266.17 116.22
CHAOS 200 0.7356 0.8205 0.6662 0.653 77.31 74.05 93.48 94.5
250 0.6982 0.7698 0.5687 0.7716 148.26 103.01 107.56 36.01
300 0.7402 0.797 0.5978 0.6667 130.12 48.79 103.09 112.75
350 0.6711 0.7872 0.5516 0.6627 53.11 48.75 101.34 99.07
0.F.2 Comparison of the Selection of Variance
Table 6: Illustration of the number of selection of variance
Dataset $\sigma^{2}$ Dice Score↑ HD95↓
Spleen Liver Left Kidney Right Kidney Spleen Liver Left Kidney Right Kidney
BTCV 0.1 0.5104 0.7478 0.5474 0.5192 394.87 299.2 329.39 273.61
1 0.8279 0.8157 0.7599 0.7164 63.09 265.79 266.17 116.22
10 0.5754 0.7691 0.6103 0.3468 388.11 312.36 238.93 323.79
CHAOS 0.1 0.5472 0.5707 0.3669 0.3813 56.26 51.06 31.72 56.14
1 0.7402 0.8205 0.6662 0.7716 53.11 48.75 93.48 36.01
10 0.5399 0.5413 0.4058 0.6664 164.89 45.93 68.55 112.75

### v2

@@ 6 Acknowledgment
This work was supported in part by U.S. NIH grants R01GM134020 and P41GM103712, NSF grants DBI-1949629, DBI-2238093, IIS-2007595, IIS-2211597, and MCB-2205148. This work was supported in part by UPMC Enterprises, Oracle Cloud credits and related resources provided by Oracle for Research, and the computational resources support from AMD HPC Fund.
Appendix A Sampling of the Dependency Scores
@@
During inference, to approximate the integral of Eq. 9, we sample all the dependency scores independently for $M$ times and calculate the final segmentation output $y^{*}$ where $\alpha_{m}$ denotes the sampled dependency scores:
@@
Given that the dependency scores within the same PMSA layer are independent of each other, and the dependency scores of deeper PMSA layer are dependent on those of former PMSA layers, as indicated by Eq. 17, Eq. 9 could be written as: